OpenStudy (anonymous):

-3+4(x-3) = (x-3) please help and show your work.

OpenStudy (anonymous):

-3+4x-12=x-3 4x-15=x-3 -x = -x --------------- 3x-15=-3 +15=+15 ----------- 3x=12 x=4 check -3+4(4-3)=4-3 -3+4(1)=1 1=1 check

OpenStudy (anonymous):

-3+4x-12=x-3 -15+4x=x-3 4x-x=15-3 3x=12 x=4

OpenStudy (saifoo.khan):

-3 +4x - 12 = x - 3 4x -15 = x - 3 4x - x = 15 -3 3x = 12 x = 12/3 x = 4